What is the gain in potential energy for a 77kg person when climbing 20 stair steps if each step is 72.4 centimeters high ? Hint: PE=m⋅g⋅h.
Answer #1

The given values of this question are as follows

Mass of person =77kg

Height of each step(h​​​​1​​​)=72.4c.m.=0.724m. [As0.01m.=1c.m.]

And also the man climbed 20steps ,So the total height is (h)=20*0.724m.=14.48m

And we know that "g" is the acceleration due to gravity and it's value is given by 9.8m/s​​​​​​​​​2

By putting all the values in the expression P.E=m*g*h we can calculate potential gain by the person.So, let's put and calculate

P.E=77*9.8*14.48

P.E=10926.608 joules

P.E=10926.608 J

Summary:-​​​​​​-So,the gain in potential energy for a 77kg person when climbing 20 stair steps if each step is 72.4 centimeters high is 10926.608J.

And here in this question 77kg person indicates that a person whose mass is 77kg not the weight ( because this is a constant and we know that mass is constant whole over the universe while weight varies from planet to planet).